nicks boat can travel 36 miles upstream in the same time that it can travel 54 miles downstream. the speed of the current is 3 mph, find the speed of nicks boat without the current.
thanks!

Answer by checkley77(12844)

36/(S-3)=54/(S+3)
54(S-3)=36(S+3)
54S-162=36S+108
54S-36S=108+162
18S=270
S=270/18
S=15 MPH IS THE SPEED OF THE BOAT IN STILL WATER.
PROOF:
36/(15-3)=54/(15+3)
36/12=54/18
3=3